Suppose you are given a one mole of ideal gas, in an initial state p=6 atm, v=3L, and in a clock wise direction, you maneuvered the system reverisbly in a path given by (v-6)2

+(p-6)2 =16 in the pv plane

Deduce the amount of work done by the gas as a result of this process

Find out the change if the internal energy of the gas

Calculate the minimum and maximum temperature attained by the gas during the cycle



1
Expert's answer
2021-05-10T09:23:09-0400

"p = 6\ atm\\
V = 3\ L"

"path = (V-6)^2\n+(p-6)^2\ufeff\ufeff =16"


maximum value of V when p = 6 atm

"(V-6)² + (6-6)² = 16\\
(V-6)² = 16\\
V-6 = 4\\
V= 10\ L"


maximum value of p when V = 3 L

"(3-6)²+ (p -6)² = 16\\
(-3)² +(p-6)² =16\\
9+ (p-6)²= 16\\
(p-6)² = 7\\
p-6 =\sqrt{7}\\
p= 6 +\sqrt7= 8.65 \ atm"


Work done = "\dfrac12(V_f-V_i)(p_f-p_i)"


"=\dfrac12(10-3)(8.65-6) = 9.275\ J"



"\u2206U = Q-W.D = 0-9.275J"

"\therefore ∆U = -9.275\ J"